The owner of a restaurant that serves continental food wants to study characteristics of his customers. He decides to focus on two variables: the amount of money spent by customers and whether customers order dessert. The results from a sample of 60 customers are as follows: • Amount spent: X=$38.54, S= $7.26 • 18 customers purchased dessert. a. Construct a 95% confidence interval estimate of the population mean amount spent per customer in the restaurant.



1
Expert's answer
2022-05-30T03:55:21-0400

Confidence interval:

"\=x\pm z\frac{s}{\sqrt{n}}"

z is the chosen Z-value (1.96 for 95% from t-table)

"38.54-(1.96)(\frac{7.26}{\sqrt{60}})<\mu<38.54+(1.96)(\frac{7.26}{\sqrt{60}})"

"36.70<\mu<40.38"

95% confidence interval is (36.70, 40.38)